OpenStudy (anonymous):

A parallelogram with the sides of 6 and 10 has a area of 30 square root of 2. Find the measure of each angle of the parallelogram. Small angle-? Large angle-?

OpenStudy (anonymous):

the area : A = 30 square root of 2 height : h A = 10h => h = A/10 = 3 square root of 2 sin(small angle) = (3 square root of 2)/6 small angle = 45º big angle = 135º

OpenStudy (anonymous):

wow thanks couldn't of done it without you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Are you fine with this answer?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i have a new question